OpenStudy (anonymous):

y = 5x and 2x + 3y = 34 (- 5x + y = 0)(-3)→ 15x - 3y = 0 2x + 3y = 34 → 2x + 3y = 34 2x + 3(5x) = 34 5x = 2x + 3y 5x + (2x + 3y) = 34 5x - (2x - 3y) = -34 its the second one right?

OpenStudy (ash2326):

I'm trying to get this part 2x + 3(5x) = 34 5x = 2x + 3y 5x + (2x + 3y) = 34 5x - (2x - 3y) = -34 What have you tried to do?

OpenStudy (anonymous):

oops sorry i missed a part of the proiblem my bad Choose the correct first step(s) to solve the system using substitution. y = 5x and 2x + 3y = 34 A) (- 5x + y = 0)(-3)→ 15x - 3y = 0 2x + 3y = 34 → 2x + 3y = 34 B) 2x + 3(5x) = 34 C) 5x = 2x + 3y D) 5x + (2x + 3y) = 34 E) 5x - (2x - 3y) = -34

OpenStudy (anonymous):

i think its B

OpenStudy (ash2326):

Yeah, it's right :)

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

im sure b4 the prob sounded pretty weird haha and hard

OpenStudy (ash2326):

yeah, I was stuck :P

OpenStudy (anonymous):

hahaha

OpenStudy (ash2326):

=D

OpenStudy (anonymous):

XD